# Reactive capacitance of flat patches of arbitrary shape

### v1

@@ III.2 Effect of patch anisotropy onto the eigenvalues
$\mu_{0}\approx\frac{1}{|\Gamma|\mathcal{A}_{\Gamma}}\simeq\frac{\pi}{a(1+2\ln(2b/a))}\qquad(a\ll b).$ (94)
@@
Figure 8: Asymptotic behavior of the principal eigenvalue $\mu_{0}$ for elliptic, rectangular and rhombic patches with $b=1$ and variable $a$. Symbols show $1/(a\mu_{0})$ as a function of $\ln(b/a)$, whereas lines indicate linear fits given in the text.
@@ III.3 Spectral weights
Figure 9: Relative contribution $F_{0}$ of the principal eigenmode as a function of $a/b$, for rectangles, ellipses and rhombuses, with $b=1$ and variable $a$.
The empirical observation that $F_{0}$ remains above $0.78$ for the considered domains and aspect ratios raises several geometric questions. Even though $F_{0}$ decreases on the range $0.01\leq a/b\leq 1$, it is yet unclear whether this decay holds in the limit $a\to 0$. Moreover, even if $F_{0}\to 0$ as $a\to 0$, such a slow decay requires to consider extremely thin patches to achieve small $F_{0}$. Is it possible to construct other types of shapes to diminish $F_{0}$? While the actual value of $F_{0}$ does not look relevant, the closeness of $F_{0}$ to $1$ may have important consequences. For instance, the upper and lower bounds (84, 85) are close when $F_{0}\approx 1$, yielding thus an accurate estimate for the principal eigenvalue $\mu_{0}$. Moreover, we showed in Sec. II.4 that the accuracy of the sigmoidal approximation (27) is higher when $F_{0}$ is close to $1$. For these reasons, further analysis and eventual decrease of $F_{0}$ may provide a better understanding of the reactive capacitance.
@@ III.4 Sigmoidal approximation
According to Eq. (34), the relative error of the sigmoidal approximation cannot exceed ${\mathcal{E}}_{\rm max}$. In the case of elliptic patches, one can use the exact relations (90, 143) for $C(\infty)$ and $\mathcal{A}_{\Gamma}$ to get ${\mathcal{E}}_{\rm max}=32/(3\pi^{2})-1\approx 8.1\%$. This conservative bound is twice larger than the actual maximal relative error. For the unit square patch, Eq. (140) yields $\mathcal{A}_{\Gamma}\approx 0.4733$, whereas the electrostatic capacitance was found numerically to be $C(\infty)\approx 0.3667874$ [104] so that ${\mathcal{E}}_{\rm max}\approx 9.1\%$. This bound is again twice larger than the actual maximal relative error.
@@ IV Discussion and conclusion
$C(\infty)=\frac{|\Gamma|}{2\pi}\biggl(\mu_{0}\underbrace{\sum\limits_{k=0}^{\infty}F_{k}}_{=1}+\sum\limits_{k=1}^{\infty}F_{k}(\mu_{k}-\mu_{0})\biggr)$ (97)
@@
Another related open question is to characterize the patches, for which $F_{0}$ is smaller than a given threshold. According to our numerical results, one may probably achieve this condition by taking extremely thin elliptic or rhombic patches, but are there other ways to diminish $F_{0}$? We provided an example of a dumbbell domain, for which the relatively low value $F_{0}\approx 0.56$ could result from the localization phenomenon. However, a rigorous analysis of this phenomenon via local estimates on the eigenfunctions seems to be much more challenging than in the conventional setting of Laplacian eigenfunctions (see [103]). Moreover, one can inspect the contribution of other eigenfunctions and ask whether there exist patches, for which the relative contribution of the principal eigenfunction is not maximal? Even though these questions sound rather abstract, a better understanding of the relation between the shape of the patch and the geometry of its Steklov eigenfunctions may help to design more efficient catalysts with desired reactive properties.
@@
The second practically important result is that the simple sigmoidal approximation $C^{\rm app}(\mu)$ from Eq. (26) turns out to be accurate for a variety of patch shapes that we considered. This numerical observation suggests that, for many practical purposes, it is sufficient to know the electrostatic capacitance $C(\infty)$ and the surface area $|\Gamma|$ of the patch. The additive form of Eq. (26) allows one to interpret $C^{\rm app}(\mu)$ as a consecutive connection of a “diffusion resistance” $1/C(\infty)$ and a “reaction resistance” $2\pi/(\mu|\Gamma|)$, in the language of electrostatics. In fact, as a resistance is the ratio between the applied voltage and the induced electric current, $1/C(\mu)$ plays the role of a resistance in diffusion-reaction problems. The above interpretation was proposed as an empirical way to describe the Laplacian transport, which is decomposed into the first diffusive step and the second reactive step [96, 97, 98, 99]. Its mathematical validation for electrochemical transport was provided in [82], while our work justifies it for steady-state diffusion-controlled reactions on partially reactive flat patches. Moreover, we obtained an upper bound (34) for the relative error of the sigmoidal approximation, which is expressed in terms of $|\Gamma|$, $C(\infty)$, and the constant $\mathcal{A}_{\Gamma}$ defined in Eq. (20). We derived a useful representation (136) for an efficient numerical computation of $\mathcal{A}_{\Gamma}$ for general polygonal patches and provided the exact explicit formulas for elliptic and rectangular patches. The versatile roles of this constant in various contexts were discussed in [81].
@@
Following [90, 81], we introduce the reactive size $L_{\Gamma}=4|\Gamma|/(\pi^{2}C(\infty))$ of a flat patch of arbitrary shape. The numerical prefactor $4/\pi^{2}$ was included for convenience, to ensure that the reactive size of a circular patch is its diameter $2a$. For instance, the reactive size of a quadratic patch with edge length $L$ is $L_{\Gamma}\approx 1.11L$, where we used the numerical value $C(\infty)\approx 0.3667874$ for the capacitance of the unit square [104]. For an elliptic patch with semiaxes $a<b$, one has
$L_{\Gamma}=\frac{4a\,K(\sqrt{1-a^{2}/b^{2}})}{\pi}\,.$ (98)
@@
For elongated elliptic patches with $a\ll b$, one gets $L_{\Gamma}\approx\tfrac{4}{\pi}a\ln(4b/a)$, i.e., the reactive size vanishes as $a\to 0$, as expected.
@@ A.1 Finite-element method
Up to now, this was a standard procedure for an arbitrary kernel $\mathcal{G}(\bm{x},\bm{y})$. In the following, we rely on the explicit form of the kernel and the flat geometry of the patch to obtain a semi-analytical representation of the matrix elements $G_{jk,j^{\prime}k^{\prime}}$. In fact, we aim to evaluate exactly the integral
@@ B.2 Rectangular patch
$\mathcal{A}_{\Gamma}=\frac{2\eta(1)}{\pi b}\,$, (140)
@@ B.4 Circular annulus
$\mathcal{A}^{io}_{\Gamma}$ $=\frac{R_{1}R_{2}\sqrt{R_{1}^{2}+R_{2}^{2}}}{3|\Gamma|^{2}}\int\limits_{0}^{2\pi}dt\biggl(\cos(t)\sqrt{1-\gamma\cos(t)}$
$+\frac{\cos(t)-\tfrac{\gamma}{2}(1+\cos^{2}(t))}{\sqrt{1-\gamma\cos(t)}}\biggr),$
@@
By symmetry, one also gets $\mathcal{A}^{oi}_{\Gamma}=\mathcal{A}^{io}_{\Gamma}$ that completes the computation of $\mathcal{A}_{\Gamma}$.

### v2

@@ III.2 Effect of patch anisotropy onto the eigenvalues
$\mu_{0}\approx\frac{1}{|\Gamma|\mathcal{A}_{\Gamma}}\simeq\frac{\pi}{a(1+2\ln(2b/a))}\qquad(a\ll b),$ (95)
@@
Figure 8: Asymptotic behavior of the principal eigenvalue $\mu_{0}$ for elliptic, rectangular and rhombic patches with $b=1$ and variable $a$. Symbols show $1/(a\mu_{0})$, obtained numerically, as a function of $\ln(b/a)$, whereas lines indicate linear fits given in the text.
@@ III.3 Spectral weights
Figure 9: Relative contribution $F_{0}$ of the principal eigenmode, obtained numerically, as a function of $a/b$, for rectangles, ellipses and rhombuses, with $b=1$ and variable $a$.
The empirical observation that $F_{0}$ remains above $0.78$ for the considered domains and aspect ratios raises several geometric questions. Even though $F_{0}$ decreases on the range $0.01\leq a/b\leq 1$, it is yet unclear whether this decay holds in the limit $a\to 0$. Moreover, even if $F_{0}\to 0$ as $a\to 0$, such a slow decay requires to consider extremely thin patches to achieve small $F_{0}$. Is it possible to construct other types of shapes to diminish $F_{0}$? While the actual value of $F_{0}$ does not look relevant, the closeness of $F_{0}$ to $1$ may have important consequences. For instance, the upper and lower bounds (85, 86) are close when $F_{0}\approx 1$, yielding thus an accurate estimate for the principal eigenvalue $\mu_{0}$. Moreover, we showed in Sec. II.4 that the accuracy of the sigmoidal approximation (27) is higher when $F_{0}$ is close to $1$. For these reasons, further analysis and eventual decrease of $F_{0}$ may provide a better understanding of the reactive capacitance.
@@ III.4 Sigmoidal approximation
According to Eq. (34), the relative error of the sigmoidal approximation cannot exceed ${\mathcal{E}}_{\rm max}$. In the case of elliptic patches, one can use the exact relations (91, 146) for $C(\infty)$ and $\mathcal{A}_{\Gamma}$ to get ${\mathcal{E}}_{\rm max}=32/(3\pi^{2})-1\approx 8.1\%$. This conservative bound is twice larger than the actual maximal relative error. For the unit square patch, Eq. (143) yields $\mathcal{A}_{\Gamma}\approx 0.4733$, whereas the electrostatic capacitance was found numerically to be $C(\infty)\approx 0.3667874$ [104] so that ${\mathcal{E}}_{\rm max}\approx 9.1\%$. This bound is again twice larger than the actual maximal relative error.
@@ IV Discussion and conclusion
$C(\infty)=\frac{|\Gamma|}{2\pi}\biggl(\mu_{0}\underbrace{\sum\limits_{k=0}^{\infty}F_{k}}_{=1}+\sum\limits_{k=1}^{\infty}F_{k}(\mu_{k}-\mu_{0})\biggr),$ (98)
@@
Another related open problem is to characterize the patches, for which $F_{0}$ is smaller than a given threshold. According to our numerical results, one may probably achieve this condition by taking extremely thin elliptic or rhombic patches, but are there other ways to diminish $F_{0}$? We provided an example of a dumbbell domain, for which the relatively low value $F_{0}\approx 0.56$ could result from the localization phenomenon. However, a rigorous analysis of this phenomenon via local estimates on the eigenfunctions seems to be much more challenging than in the conventional setting of Laplacian eigenfunctions (see [103]). Moreover, one can inspect the contribution of other eigenfunctions and ask whether there exist patches, for which the relative contribution of the principal eigenfunction is not maximal? Even though these questions sound rather abstract, a better understanding of the relation between the shape of the patch and the geometry of its Steklov eigenfunctions may help to design more efficient catalysts with desired reactive properties.
@@
The second practically important result is that the simple sigmoidal approximation $C^{\rm app}(\mu)$ from Eq. (26) turns out to be accurate for a variety of patch shapes that we considered. This numerical observation suggests that, for many practical purposes, it is sufficient to know the electrostatic capacitance $C(\infty)$ and the surface area $|\Gamma|$ of the patch. The additive form of Eq. (26) allows one to interpret $C^{\rm app}(\mu)$ as a consecutive connection of a “diffusion resistance” $1/C(\infty)$ and a “reaction resistance” $2\pi/(\mu|\Gamma|)$, in the language of electrostatics. In fact, as a resistance is the ratio between the applied voltage and the induced electric current, $1/C(\mu)$ plays the role of a resistance in diffusion-reaction problems. The above interpretation was proposed as an empirical way to describe the Laplacian transport, which is decomposed into the first diffusive step and the second reactive step [96, 97, 98, 99]. Its mathematical validation for electrochemical transport was provided in [82], while our work justifies it for steady-state diffusion-controlled reactions on partially reactive flat patches. Moreover, we obtained an upper bound (34) for the relative error of the sigmoidal approximation, which is expressed in terms of $|\Gamma|$, $C(\infty)$, and the constant $\mathcal{A}_{\Gamma}$ defined in Eq. (20). We derived a useful representation (139) for an efficient numerical computation of $\mathcal{A}_{\Gamma}$ for general polygonal patches and provided the exact explicit formulas for elliptic and rectangular patches. The versatile roles of the constant $\mathcal{A}_{\Gamma}$ in various contexts were discussed in [81].
The reactive capacitance was shown to determine the small-patch asymptotic behavior of many characteristics of diffusion-controlled reactions, such as the mean first-passage time, splitting probabilities, and the total flux [68, 69]. For instance, for restricted diffusion inside a bounded domain $\Omega$ with a reflecting boundary covered by multiple small reactive patches, well-separated from each other, the leading-order term of the volume-averaged mean first-reaction time is
$T\sim\frac{|\Omega|}{2\pi D\overline{C}}\,$, (99)
where $\overline{C}$ is the sum of reactive capacitances of the patches, $D$ is the diffusivity, and $|\Omega|$ is the volume of the domain. One sees that the reactive capacitance plays the role of an effective patch size that incorporates its shape and reactivity. For a single patch, our sigmoidal approximation (26) yields
$T\sim|\Omega|\biggl(\frac{1}{2\pi DC(\infty)}+\frac{1}{\kappa|\Gamma|}\biggr),$ (100)
in agreement with the above discussion on the consecutive connection of the diffusive and reactive steps. Two next-order correction terms were derived in [68].
@@
Following [90, 81], we also introduce the reactive length $L_{\Gamma}=|\Gamma|/(2\pi C(\infty))$, associated to a flat patch of arbitrary shape, in order to rewrite the approximate reactive capacitance as $C^{\rm app}(\mu)=C(\infty)/(1+1/(\mu L_{\Gamma}))$. This expression indicates how the reactivity of the patch reduces its electrostatic capacitance $C(\infty)$, and $L_{\Gamma}$ is the proper geometric lengthscale that is compared to the physical lengthscale $1/\mu=D/\kappa$ (see [96, 97] for further discussions). For instance, the reactive length of a circular patch of radius $a$ is $\tfrac{\pi}{4}a$, whereas the reactive length of a quadratic patch with edge length $L$ is $L_{\Gamma}\approx 0.43\,L$ (here we used the numerical value $C(\infty)\approx 0.3667874$ for the capacitance of the unit square [104]). For an elliptic patch with semiaxes $a<b$, we get
$L_{\Gamma}=a\frac{K(\sqrt{1-a^{2}/b^{2}})}{2}\,.$ (101)
@@
When the patch is elongated ($a\ll b$), one finds $L_{\Gamma}\approx\tfrac{1}{2}a\ln(4b/a)$, i.e., the reactive length vanishes as $a\to 0$, as expected.
@@ A.1 Finite-element method
Up to now, this was a standard numerical procedure for an arbitrary kernel $\mathcal{G}(\bm{x},\bm{y})$. In the following, we rely on the explicit form of the kernel and the flat geometry of the patch to obtain a semi-analytical representation of the matrix elements $G_{jk,j^{\prime}k^{\prime}}$. In fact, we aim to evaluate exactly the integral
@@ B.2 Rectangular patch
$\mathcal{A}_{\Gamma}=\frac{2\eta(1)}{\pi b}\approx\frac{0.2366}{b}\,$, (143)
@@ B.4 Circular annulus
$\mathcal{A}^{io}_{\Gamma}$ $=\frac{R_{1}R_{2}\sqrt{R_{1}^{2}+R_{2}^{2}}}{3|\Gamma|^{2}}\int\limits_{0}^{2\pi}\biggl(\cos(t)\sqrt{1-\gamma\cos(t)}$
$+\frac{\cos(t)-\tfrac{\gamma}{2}(1+\cos^{2}(t))}{\sqrt{1-\gamma\cos(t)}}\biggr)dt,$
@@
Using the Landen’s transformations, one gets
$K(\sqrt{2\gamma/(1+\gamma)})=\frac{2K((1-\sqrt{1-\gamma^{2}})/\gamma)}{1+\sqrt{(1-\gamma)/(1+\gamma)}}$
$\quad=(1+R_{1}/R_{2})K(R_{1}/R_{2}),$
$E(\sqrt{2\gamma/(1+\gamma)})=\biggl(1+\sqrt{\frac{1-\gamma}{1+\gamma}}\biggr)E((1-\sqrt{1-\gamma^{2}})/\gamma)$
$\quad-\frac{2}{1+\sqrt{\frac{1+\gamma}{1-\gamma}}}K((1-\sqrt{1-\gamma^{2}})/\gamma)$
$\quad=\frac{2E(R_{1}/R_{2})}{1+R_{1}/R_{2}}-(1-R_{1}/R_{2})K(R_{1}/R_{2}).$
Substituting these representations, we get after simplifications:
$\mathcal{A}^{io}_{\Gamma}=\frac{4R_{2}\bigl[(R_{2}^{2}-R_{1}^{2})K(R_{1}/R_{2})-(R_{2}^{2}+R_{1}^{2})E(R_{1}/R_{2})\bigr]}{3|\Gamma|^{2}}.$
By symmetry, one also gets $\mathcal{A}^{oi}_{\Gamma}=\mathcal{A}^{io}_{\Gamma}$ that completes the computation of $\mathcal{A}_{\Gamma}$:
$\mathcal{A}_{\Gamma}$ $=\frac{8}{3\pi^{2}(R_{2}^{2}-R_{1}^{2})^{2}}\biggl(R_{1}^{3}+R_{2}^{3}+R_{2}(R_{2}^{2}-R_{1}^{2})K(R_{1}/R_{2})$
$-R_{2}(R_{2}^{2}+R_{1}^{2})E(R_{1}/R_{2})\biggr).$ (152)
In contrast, there is no an explicit exact representation of the electrostatic capacitance $C(\infty)$ of a circular annulus (see the discussion in Sec. 8.6 of [106]); an approximate formula was given in [107].